OpenStudy (anonymous):

Given the function f(x) = 4x - 16, what is the x-value of the locator point of f-1(x)?

OpenStudy (campbell_st):

is the locator point, where the curves f(x) and the inverse intersect...?

OpenStudy (anonymous):

don't really know

OpenStudy (korosh23):

Do it like this. Y= 4x-16 16=4x 16/4=x 4=x value of x

OpenStudy (korosh23):

Is this what you are looking for ?

OpenStudy (anonymous):

is that how you are suppose to do the problem?

OpenStudy (anonymous):

its suppose to be 4^x not 4x, I messed that up

OpenStudy (korosh23):

Is it possible if you rewrite your equation?

OpenStudy (anonymous):

Given the function f(x) = 4^x - 16, what is the x-value of the locator point of f^-1 (x)?

OpenStudy (campbell_st):

well the inverse function is \[f(x) = \frac{ x + 16}{4}\] ok, doing some reading the locator point for a linear function is the y- intercept... so the y- intercept has coordinates (0, a) where y = a is the intercept... so the x value of the inverse is x = 0 that's my best guess here is a clipping of some notes I found

OpenStudy (korosh23):

It is kind of based on common sense.

OpenStudy (campbell_st):

so find the inverse of \[y = 4^x - 16\]

OpenStudy (korosh23):

Y= 4^x -16 16= 4^x

OpenStudy (korosh23):

In here the value of x is 2

OpenStudy (korosh23):

4^2=16

OpenStudy (korosh23):

I hope I answered your question.

OpenStudy (anonymous):

oh ok thank you for making the problem clear to understand

OpenStudy (korosh23):

No problem. :)